Determine the annual equivalent amount over 4 years for the following cash flow (CF) pattern. The nominal annual interest rate is 3%, and it is compounded weekly

EOY

0

1

2

3

4

CF

-$7500

-$2000

$14000

$18000

$4000
Answer #1

Assuming 1 yr = 52 weeks

Effective interest rate = (1+0.03 / 52)^52 - 1

= (1+0.000576923)^52 - 1

= (1.000576923)^52 - 1

= 0.0304456

NPW of cash flow = -7500 - 2000/ (1+0.0304456) + 14000 / (1+0.0304456)^2 + 18000 / (1+0.0304456)^3 + 4000 / (1+0.0304456)^4

= -7500 - 2000/ (1.0304456) + 14000 / (1.0304456)^2 + 18000 / (1.0304456)^3 + 4000 / (1.0304456)^4

= 23743.02

Equivalent annual worth = 23743.02 * (A/P,3.04456%,4)

= 23743.02 * 0.0304456 * ((1 + 0.0304456)^4)/((1 + 0.0304456)^4-1)

= 23743.02 * 0.0304456 * ((1.0304456)^4)/((1.0304456)^4-1)

= 6394.32